OpenStudy (anonymous):

What is the solution to the rational equation 3 x over 4 minus the fraction 1 over 4 = 5 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\Large \frac{ 3x }{ 4 }-\frac{ 1 }{ 4 }=5\] Add 1/4 to both sides of the equation \[\Large \frac{ 3x }{ 4 }=5+\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

simplify that by adding the 5 to 1/4... and what do you get??

OpenStudy (anonymous):

the answers are 5,4,7,or 9

OpenStudy (anonymous):

3x/4 = 5.25